# SHINE: Saliency-aware HIerarchical NEgative Ranking for Compositional Temporal Grounding

### v1

@@ Abstract
Temporal grounding, a.k.a video moment retrieval, aims at locating video segments corresponding to a given query sentence. The compositional nature of natural language enables the localization beyond predefined events, posing a certain challenge to the compositional generalizability of existing methods. Recent studies establish the correspondence between videos and queries through a decompose-reconstruct manner to achieve compositional generalization. However, they only consider dominant primitives and build negative queries through random sampling and recombination, resulting in semantically implausible negatives that hinder the models from learning rational compositions. In addition, recent DETR-based methods still underperform in compositional temporal grounding, showing irrational saliency responses when given negative queries that have subtle differences from positive queries. To address these limitations, we first propose a large language model-driven method for negative query construction, utilizing GPT-3.5-Turbo to generate semantically plausible hard negative queries. Subsequently, we introduce a coarse-to-fine saliency ranking strategy, which encourages the model to learn the multi-granularity semantic relationships between videos and hierarchical negative queries to boost compositional generalization. Extensive experiments on two challenging benchmarks validate the effectiveness and generalizability of our proposed method. Our code is available at https://github.com/zxccade/SHINE.
@@ 1 Introduction
Temporal grounding [8, 17, 52, 50, 18, 29] in videos has received continuous attention for its wide range of applications in moment retrieval and automatic content generation. Unlike typical temporal action localization tasks [54], temporal grounding aims to retrieve video moments based on textual queries that include not only the action itself but also the objects, attributes, and interactions involved. Moreover, the extensive range of vocabulary in natural language can expand the limited words from labeled videos to describe a variety of new things and scenarios unlabeled in training data. However, recent works [47, 5, 34, 33] have shown that existing vision-language models (VLMs) lack compositional generalizability for unseen combinations, as reflected by their insensitivity to word order and significant primitives.
@@ 2 Related Work
Temporal Grounding. Temporal grounding, a.k.a. video moment retrieval, initially proposed in [8] and [17], aims at localizing segments in a video that match the description of a query sentence. Currently, dominant supervised learning techniques are divided into two categories: proposal-based and proposal-free methods. Proposal-based methods generate candidate segments through various strategies, including sliding windows[8, 24], dense proposals[40, 37, 12], and fixed anchors[4, 49, 45, 52, 35], subsequently selecting the most appropriate intervals based on a similarity measure. However, the generation of candidate proposals and their semantic matching with queries are computationally intensive. Conversely, proposal-free methods [46, 30, 48, 11, 41, 22] directly predict the temporal boundaries of the target clip. This paradigm eliminates the need for proposal generation, significantly enhancing the model’s efficiency during inference. Recently, Lei et al. [18] reformulated the temporal grounding task as a set prediction problem, introducing a DETR-based [2] architecture enabling simultaneous video moment retrieval and highlight detection. Subsequent works, including UMT [25], QD-DETR [29], and EaTR [14], have enhanced localization accuracy by refining the DETR framework. Differently, our work establishes a connection between the saliency score and compositionality, which effectively unlocks the potential for compositional generalization in DETR-based models.
@@ 3.1 Problem Definition and Overview
The overall framework of our proposed method is shown in Fig. 2. Given a video-query pair $(V_{p},Q_{p})$, we first construct a set of hierarchical hard negative queries $\{Q^{i}_{hn}\}^{3}_{i=1}$ via a progressive mask-and-predict strategy (Sec. 3.2). Notably, we utilize a large language model, namely GPT-3.5-Turbo [1], to select appropriate words from the training set for primitive replacement. This operation gradually changes the semantics of the original query while effectively avoiding implausible combinations. These manipulated queries, along with the original query $Q_{p}$ and a negative query $Q_{n}$ selected from other queries in the same mini-batch, constitute a set of queries. During the training phase, each video and its associated query set are first fed into a video encoder $\mathcal{F}_{v}(\cdot)$ and a text encoder $\mathcal{F}_{t}(\cdot)$ to extract corresponding features, which are then processed by the encoder to predict the saliency scores $\{S_{p},S^{1}_{hn},S^{2}_{hn},S^{3}_{hn},S_{n}\}$. Subsequently, we propose modeling the video-level saliency prior using a coarse-grained saliency ranking loss $\mathcal{L}_{cr}$, which incorporates two constraints designed to enhance the discriminability between positive and negative queries (Sec. 3.3). Concurrently, we employ a fine-grained saliency ranking loss $\mathcal{L}_{fr}$ to discern the saliency scores derived from the query set, which facilitates the learning of multi-granularity semantics by exploring the nuance among hierarchical negative samples. By combining $\mathcal{L}_{base}$ with the proposed coarse-to-fine saliency ranking loss (Sec. 3.4), our method can be seamlessly integrated into existing DETR-based models [18, 29], greatly enhancing their potential for compositional generalization.
@@ 3.2 Hierarchical Hard Negatives Construction
Fig. 3 shows the pipeline for constructing hierarchical negative queries. Specifically, we first use spaCy111spaCy: https://spacy.io/ to perform part-of-speech (POS) tagging on all query sentences in the training set, and construct a dictionary $D$ by counting words from five types of primitives (verbs, nouns, adjectives, prepositions, and adverbs). Subsequently, for a given query $Q$, we progressively mask the primitives with different ratios in the original query according to their relative importance in linguistics, i.e., verb-noun-adjective-preposition-adverb. Also, to ensure contextual consistency in negative queries, subjects (usually nouns) are only considered when other primitives are insufficient. Instead of filling in the masks with random selections from $D$, we resort to a powerful LLM, i.e., GPT-3.5-Turbo, to generate semantically plausible hard negative queries. Considering the token limit inherent to LLMs, we randomly select a subset from the dictionary to ensure a balance between context and diversity for each sample. Furthermore, we have carefully crafted a prompt template to steer the LLM toward producing challenging negative queries. These negatives are semantically viable yet distinct from the original query, laying the foundation for fine-grained saliency ranking.
@@ 3.3 Coarse-to-Fine Saliency Ranking
Coarse-Grained Saliency Ranking. As shown in Fig. 4, for a given video-text pair $(V_{p},Q_{p})$, we designate video clips within and outside the ground-truth interval as $V_{p}^{+}$ and $V_{p}^{-}$, respectively, and $Q_{p}$ as the corresponding positive query. Concurrently, a negative query is randomly selected from the same mini-batch, denoted as $Q_{n}$. Intuitively, the saliency scores within the ground-truth interval should surpass those outside, and scores elicited by positive queries should exceed those by negative ones. With these two priors, we introduce a coarse-grained ranking loss with dual constraints, formulated as:
@@ 3.4 Model Training Objectives
where $\mathcal{L}_{base}$ represents the basic loss of the DETR-based model, typically including bipartite matching loss, moment localization loss, and saliency loss. $\alpha$ and $\beta$ are two coefficients that control the loss weights. By combining the two constraints with the basic loss function, our method can significantly enhance the model’s compositional generalization capabilities while preserving accuracy for in-distribution samples, as demonstrated in subsequent experiments.
@@ 4 Experiments
Datasets and Evaluation Metrics. We evaluate our method on two newly proposed benchmarks, Charades-CG and ActivityNet-CG[20], originated from Charades-STA [8] and ActivityNet Captions [17]. Each dataset is reorganized into four splits: Training/Test-Trivial/Novel-Composition/Novel-Word, where the latter three splits evaluate the model’s performance on IID samples, novel combinations of seen words, and unseen words, respectively. In particular, the Novel-Composition split considers five types of new compositions: verb-noun, noun-noun, verb-adverb, adjective-noun, and preposition-noun. Following previous works [20, 19, 42], we use two main metrics to evaluate our methods, i.e., "$\text{R}@\text{n},\text{IoU}=m$" and mean Intersection over Union (mIoU).
Implementation Details. We adopt Moment-DETR[18] and QD-DETR[29] as our baselines and integrate them with our method using their officially released code. Unless specifically noted, other hyperparameters follow their default settings. Following [42, 19], for the Charades-CG dataset, we use pretrained I3D Network [3] to extract 1024-dimensional features for each 0.5-second clip. For ActivityNet-CG, we employ pretrained C3D network [16] to extract 4096-dimensional features for 1-second video clips. For text features, we follow [18, 29] to extract CLIP [31] features with 512 dimensions for each query. In hierarchical hard negative construction, we progressively mask the original query in Charades-CG at ratios of 25%, 50%, and 75%, while the masking ratios for ActivityNet-CG are set to 10%, 30%, and 50%. For Charades-CG, we set the learning rates for QD-DETR and Moment-DETR to 0.0001 and 0.0002, respectively, while for ActivityNet-CG, the learning rates for both models are set to 0.0002. The coarse-grained margins $h_{1}$ and $h_{2}$ are set to 1.0 and 2.0, respectively, while the relative thresholds $m_{0}$ to $m_{3}$ are set to 0.25. The factor $q$ in Eq. 2 are set to 8 for both Charades-CG and ActivityNet-CG. All experiments are run on a single NVIDIA A100 GPU with a batch size of 32 training for 200 epochs.
@@ 4.1 Comparisons with the State-of-the-arts
Tab. 1 shows the overall performance of our approach on the Charades-CG dataset. We observe that: (1) While the baseline QD-DETR [29] outperforms the latest state-of-the-art methods (e.g., SSL [19] and DeCo[42]) in the Test-Trivial spilt, there is still a performance gap in the Novel-Composition and Novel-Word splits, indicating that it has poor compositional generalization capabilities. (2) Our method can significantly improve the compositional generalizability of QD-DETR, elevating 7.93% and 6.60% in R1@0.5 and R1@0.7 in Novel-Composition split, respectively, finally outperforming on all three test splits. (3) Our method can be integrated into existing DETR-based models to unlock their compositional generalizability. In Tab. 2, we also achieve competitive results on the ActivityNet-CG dataset across two baselines. In the Novel-Composition split, our method promotes the performance of QD-DETR by 2.65%, 3.43%, and 1.43% in R1@0.5, R1@0.7 and mIoU, respectively. Additionally, by integrating our approach, Moment-DETR’s performance experiences a boost of 1.39%, 0.69%, and 1.50% in R1@0.5, R1@0.7 and mIoU metrics, respectively. Notably, VISA [20] leverages external knowledge from off-the-shelf object detectors and action recognition models while our method is conducted in an end-to-end manner. Although our method underperforms VISA [20] in the Test-Trivial split, it still achieves comparable performance in the Novel-Composition split, which proves that our method has a better capability of compositional generalization.
@@ 4.2 Ablation Studies
Comparison of Hierarchical Negative Queries. We compare the effects of the random sample-based and LLM-based hard negative queries in Tab. 5. From the Charades-CG results, we observe that LLM-based negative queries significantly outperform random sampling, with improvements of 2.82% on R1@0.5 and 1.64% on mIoU. Among all the evaluated LLMs, the negative samples generated by GPT-3.5 Turbo perform better. We also note that the improvement using the LLM-based approach over random sampling is less significant in ActivityNet-CG. A possible reason is that the query sentences are longer while the replacement ratio is lower, so the retained context is still sufficient for localization. However, the LLM-based approach consistently outperforms random sampling in R1@0.7, indicating its potential for precise grounding.
@@ 4.3 Qualitative Analysis
We visualize the saliency scores of several cases in Charades-CG in Fig. 6 and observe that the existing work struggles with hard negative queries, showing irrational saliency responses. For instance, in Fig. 6(a), the hard negative query “Person puts the coffee cup under a table.” is even more salient than the positive query “Person puts the coffee cup in a box”, leading to imprecise moment localization. In contrast, our approach consistently improves the model’s ability to distinguish different words between positive and hard negative queries and yield hierarchical responses, thereby achieving better moment localization and compositional generalization. Additionally, our method can also catch the nuanced variation of adverbs and prepositions, such as different prepositions “into” and “behind” in Fig. 6(b). This indicates that ours are more sensitive to the semantic changes of different non-dominant primitives.
Fig. 7 illustrates several qualitative examples in three splits of Charades-CG. In the Test-Trivial split, although the queries don’t contain unseen compositions and words, our method demonstrates more precise alignment than the baseline. When encountering the Novel-composition "turns the light off" and “put a box on” and Novel-Word “corner” and “hands” in the queries, our method can still generalise well to them. The presented results indicate that our method effectively guides DETR-based models in utilizing hierarchical negative samples to enhance the generalizability of unseen compositions and unseen words.

### v2

@@ Abstract
Temporal grounding, also known as video moment retrieval, aims at locating video segments corresponding to a given query sentence. The compositional nature of natural language enables the localization beyond predefined events, posing a certain challenge to the compositional generalizability of existing methods. Recent studies establish the correspondence between videos and queries through a decompose-reconstruct manner to achieve compositional generalization. However, they only consider dominant primitives and build negative queries through random sampling and recombination, resulting in semantically implausible negatives that hinder the models from learning rational compositions. In addition, recent DETR-based methods still underperform in compositional temporal grounding, showing irrational saliency responses when given negative queries that have subtle differences from positive queries. To address these limitations, we first propose a large language model-driven method for negative query construction, utilizing GPT-3.5 Turbo to generate semantically plausible hard negative queries. Subsequently, we introduce a coarse-to-fine saliency ranking strategy, which encourages the model to learn the multi-granularity semantic relationships between videos and hierarchical negative queries to boost compositional generalization. Extensive experiments on two challenging benchmarks validate the effectiveness and generalizability of our proposed method. Our code is available at https://github.com/zxccade/SHINE.
@@ 1 Introduction
Temporal grounding [9, 19, 56, 32] has received continuous attention for its wide range of applications in video summarization [22, 13], moment retrieval [53, 25], surveillance and security [34, 35]. Unlike typical temporal action localization tasks [60], temporal grounding aims to retrieve video moments based on textual queries that include not only the action itself but also the objects, attributes, and interactions. Moreover, the extensive vocabulary of natural language can expand the limited terms from labeled videos to describe new, unlabeled scenarios in the training data. However, recent works [52, 5, 39, 38] have shown that existing vision-language models (VLMs) lack compositional generalizability for unseen combinations, as reflected by their insensitivity to word order and primitives.
@@ 2 Related Work
Temporal Grounding. Temporal grounding, a.k.a. video moment retrieval, initially proposed in [9] and [19], aims at localizing segments in a video that match the description of a query sentence. Currently, dominant supervised learning techniques are divided into two categories: proposal-based and proposal-free methods. Proposal-based methods generate candidate segments through various strategies, including sliding windows[9, 27], dense proposals[45, 42, 14], and fixed anchors[4, 55, 58, 40], subsequently selecting the most appropriate intervals based on a similarity measure. However, the generation of candidate proposals and their semantic matching with queries are computationally intensive. Conversely, proposal-free methods [51, 33, 54, 47, 25] directly predict the temporal boundaries of the target clip. This paradigm eliminates the need for proposal generation, significantly enhancing the model’s efficiency during inference. Recently, Lei et al. [20] reformulated the temporal grounding task as a set prediction problem, introducing a DETR-based [2] architecture enabling simultaneous video moment retrieval and highlight detection. Subsequent works, including UMT [28], QD-DETR [32], and EaTR [16], have enhanced localization accuracy by refining the DETR framework. Differently, our work establishes a connection between the saliency score and compositionality, which effectively unlocks the potential for compositional generalization in DETR-based models.
@@ 3.1 Problem Definition and Overview
The overall framework of our proposed method is shown in Fig. 2. Given a video-query pair $(V_{p},Q_{p})$, we first construct a set of hierarchical hard negative queries $\{Q^{i}_{hn}\}^{3}_{i=1}$ via a progressive mask-and-predict strategy (Sec. 3.2). Notably, we utilize a large language model, namely GPT-3.5 Turbo [1], to select appropriate words from the training set for primitive replacement. This operation gradually changes the semantics of the original query while effectively avoiding implausible combinations. These manipulated queries, along with the original query $Q_{p}$ and a negative query $Q_{n}$ selected from other queries in the same mini-batch, constitute a set of queries. During the training phase, each video and its associated query set are first fed into a video encoder $\mathcal{F}_{v}(\cdot)$ and a text encoder $\mathcal{F}_{t}(\cdot)$ to extract corresponding features, which are then processed by the encoder to predict the saliency scores $\{S_{p},S^{1}_{hn},S^{2}_{hn},S^{3}_{hn},S_{n}\}$. Subsequently, we propose modeling the video-level saliency prior using a coarse-grained saliency ranking loss $\mathcal{L}_{cr}$, which incorporates two constraints designed to enhance the discriminability between positive and negative queries (Sec. 3.3). Concurrently, we employ a fine-grained saliency ranking loss $\mathcal{L}_{fr}$ to discern the saliency scores derived from the query set, which facilitates the learning of multi-granularity semantics by exploring the nuance among hierarchical negative samples. By combining $\mathcal{L}_{base}$ with the proposed coarse-to-fine saliency ranking loss (Sec. 3.4), our method can be seamlessly integrated into existing DETR-based models [20, 32], greatly enhancing their potential for compositional generalization.
@@ 3.2 Hierarchical Hard Negatives Construction
Fig. 3 shows the pipeline for constructing hierarchical negative queries. Specifically, we first use spaCy111spaCy: https://spacy.io/ to perform part-of-speech (POS) tagging on all query sentences in the training set, and construct a dictionary $D$ by counting words from five types of primitives (verbs, nouns, adjectives, prepositions, and adverbs). Subsequently, for a given query $Q$, we progressively mask the primitives with different ratios in the original query according to their relative importance in linguistics, i.e., verb-noun-adjective-preposition-adverb. Also, to ensure contextual consistency in negative queries, subjects (usually nouns) are only considered when other primitives are insufficient. Instead of filling in the masks with random selections from $D$, we resort to a powerful LLM, i.e., GPT-3.5 Turbo, to generate semantically plausible hard negative queries. Considering the token limit inherent to LLMs, we randomly select a subset from the dictionary to ensure a balance between context and diversity for each sample. Furthermore, we have carefully crafted a prompt template to steer the LLM toward producing challenging negative queries. These negatives are semantically viable yet distinct from the original query, laying the foundation for fine-grained saliency ranking.
@@ 3.3 Coarse-to-Fine Saliency Ranking
Coarse-Grained Saliency Ranking. As shown in Fig. 4, for a given video-text pair $(V_{p},Q_{p})$, we designate video clips within and outside the ground-truth interval as $V_{p}^{+}$ and $V_{p}^{-}$, respectively, and $Q_{p}$ as the corresponding positive query. Concurrently, a negative query is randomly selected from the same mini-batch, denoted as $Q_{n}$.
Intuitively, the saliency scores within the ground-truth interval should surpass those outside, and scores elicited by positive queries should exceed those by negative ones. With these two priors, we introduce a coarse-grained ranking loss with dual constraints, formulated as:
@@ 3.4 Model Training Objectives
where $\mathcal{L}_{base}$ represents the basic loss of the DETR-based model, typically including bipartite matching loss, moment localization loss, and saliency loss. More details can be found in Appendix 0.A. $\alpha$ and $\beta$ are two weight coefficients. By combining the two constraints with $\mathcal{L}_{base}$, our method can significantly enhance the model’s compositional generalization capabilities while preserving accuracy for in-distribution samples, as demonstrated in subsequent experiments.
@@ 4 Experiments
Datasets and Evaluation Metrics. We evaluate our method on two newly proposed benchmarks, Charades-CG and ActivityNet-CG[23], originated from Charades-STA [9] and ActivityNet Captions [19]. Each dataset is reorganized into four splits: Training/Test-Trivial/Novel-Composition/Novel-Word, where the latter three splits evaluate the model’s performance on IID samples, novel combinations of seen words, and unseen words, respectively. In particular, the Novel-Composition split considers five types of new compositions: verb-noun, noun-noun, verb-adverb, adjective-noun, and preposition-noun. Following previous works [23, 21, 48], we use two main metrics to evaluate our methods, i.e., "$\text{R}@\text{n},\text{IoU}=m$" and mean Intersection over Union (mIoU). The former denotes the percentage of queries with at least one prediction whose IoU score is larger than $m$ within the top-$n$ predictions, while the latter refers to the average of IoU scores across all queries.
Implementation Details. We adopt Moment-DETR[20] and QD-DETR[32] as our baselines and integrate them with our method using their officially released code. Unless specifically noted, other hyperparameters follow their default settings. For QD-DETR, following [48, 21], we utilize pretrained I3D [3] features on the Charades-CG dataset and C3D [18] features on the ActivityNet-CG dataset, respectively. Following [20], we employ SlowFast [7] and CLIP [36] to extract hybrid visual features for Moment-DETR. As for text features, we follow [20, 32] to extract CLIP features with 512 dimensions for each query. In hierarchical hard negative construction, we progressively mask the original query in Charades-CG at ratios of 25%, 50%, and 75%, while the masking ratios for ActivityNet-CG are set to 10%, 30%, and 50%. For Charades-CG, we set the learning rates for QD-DETR and Moment-DETR to 0.0001 and 0.0002, respectively, while for ActivityNet-CG, the learning rates for both models are set to 0.0002. The coarse-grained margins $h_{1}$ and $h_{2}$ are set to 1.0 and 2.0, respectively, while the relative thresholds $m_{0}$ to $m_{3}$ are set to 0.25. The factor $q$ in Eq. 2 are set to 8 for both Charades-CG and ActivityNet-CG. All experiments are run on a single NVIDIA A100 GPU with a batch size of 32 training for 200 epochs.
@@ 4.1 Comparisons with the State-of-the-arts
Tab. 1 shows the overall performance of our approach on the Charades-CG dataset. We observe that: (1) While the baseline QD-DETR [32] outperforms the latest state-of-the-art methods (e.g., SSL [21] and DeCo[48]) in the Test-Trivial spilt, there is still a performance gap in the Novel-Composition and Novel-Word splits, indicating that it has poor compositional generalization capabilities. (2) Our method can significantly improve the compositional generalizability of QD-DETR, elevating 7.93% and 6.60% in R1@0.5 and R1@0.7 in Novel-Composition split, respectively, finally outperforming on all three test splits. (3) Our method can be integrated into existing DETR-based models to unlock their compositional generalizability. e.g., our approach notably enhances Moment-DETR and improves its performance by 7.66% and 5.23% in R1@0.5 in Test-Trivial and Novel-Composition splits, respectively.
In Tab. 2, we also achieve competitive results on the ActivityNet-CG dataset across two baselines. In the Novel-Composition split, our method promotes the performance of QD-DETR by 2.65%, 3.43%, and 1.43% in R1@0.5, R1@0.7 and mIoU, respectively. Additionally, by integrating our approach, Moment-DETR’s performance experiences a boost of 1.39%, 0.69%, and 1.50% in R1@0.5, R1@0.7 and mIoU metrics, respectively. Notably, VISA [23] leverages external knowledge from off-the-shelf object detectors and action recognition models while our method is conducted in an end-to-end manner. Although our method underperforms VISA [23] in the Test-Trivial split, it still achieves comparable performance in the Novel-Composition split, which proves that our method has a better capability of compositional generalization.
@@ 4.2 Ablation Studies
Comparison of Hierarchical Negative Queries. We compare the effects of the random sample-based and LLM-based hard negative queries in Tab. 5. From the Charades-CG results, we observe that LLM-based negative queries significantly outperform random sampling, with improvements of 2.82% on R1@0.5 and 1.64% on mIoU.
Among all the evaluated LLMs, the negative samples generated by GPT-3.5 Turbo perform better. We also note that the improvement using the LLM-based approach over random sampling is less significant in ActivityNet-CG. A possible reason is that the query sentences are longer while the replacement ratio is lower, so the retained context is still sufficient for localization. However, the LLM-based approach consistently outperforms random sampling in R1@0.7, indicating its potential for precise grounding.
@@ 4.3 Qualitative Analysis
We visualize the saliency scores of several cases in Charades-CG in Fig. 6 and observe that the existing work struggles with hard negative queries, showing irrational saliency responses. For instance, in Fig. 6(a), the hard negative query “Person takes the shoe in the closet.” is even more salient than the positive query “Person puts the towel in the closet”, leading to imprecise moment localization. In contrast, our approach consistently improves the model’s ability to distinguish different words between positive and hard negative queries and yield hierarchical responses, thereby achieving better moment localization and compositional generalization. Additionally, our method can also catch the nuanced variation of adverbs and prepositions, such as different prepositions “into” and “behind” in Fig. 6(b). This indicates that ours are more sensitive to the semantic changes of different non-dominant primitives.
Fig. 7 illustrates several qualitative examples in three splits of Charades-CG. In the Test-Trivial split, although the queries don’t contain unseen compositions and words, our method demonstrates more precise alignment than the baseline. When encountering the Novel-composition "turns the light off" and “put a box on” and Novel-Word “corner and hands” in the queries, our method can still generalise well to them. The presented results indicate that our method effectively guides DETR-based models in utilizing hierarchical negative samples to enhance the generalizability of unseen compositions and unseen words.